TUTOR SWAPS I. An asset manager who manages a fund called Veriane Fund follows a passive investment strategy and his portfolio tracks the S&P 500 Total Retums Index. The asset manager ty swap contract with a counterparty Morgan Stanley with the following terms Notional Principal: $200 million Verizon Fund pays: Total retuns on the S&P 500 Index Morgan Stanley pays: Fixed 6% Payments to be made at the end of every six months, that is, 30th June and 31st December The swap has a maturity of 1 year The return of the index is given in the below table Date 1* Jan 30 June 31 Dec S&P Index level 1800 1890 1760 Calculate the total payoff between Verizon Lad and Morgan Stanley on 1 Jan, 3o June and 31 Dedc
Answer #1

A swap contract as the name suggests is exchanging one return over another. Here, Morgan Stanley agrees to pay 6% fixed rate of interest on the notional principal in exchange for the returns on the S&P 500 Index.

On 1st January, both the parties have entered into a contract and hence no payoff will be made on that day. The value of swap contract on the initiation day is 0.

On 30th June, Morgan Stanley will pay 6% interest for 6 months on $200 million = (200 million *6% *6/12) = 6 million

and Verizon Fund will pay the change in S&P index level = (1890-1800) * 200 million/1800 = 10 million.

Net payoff = Verizon will pay $4 million to Morgan Stanley.

On 31st December, Morgan Stanley will again pay 6% interest for 6 months on $200 million = (200 million *6% *6/12) = 6 million

and Verizon Fund will pay the change in S&P index level = (1760-1890) * 200million/1800 = -14.44 million.

Ney payoff = Morgan Stanley will pay $20.44 million to Verizon.

As a part of this swap contract, Verizon has hedged itself with the risk of fluctuating S&P Index for a fixed return of 6%.

Note:-

200 Million/1800 is the number of shares that can be bought on 1st Jan with the notional principal of $200 million.